OpenStudy (anonymous):

$300 at 5% annual interest for 7 years A. $214.29 B. $11.67 C. $420.00 D. $105.00

OpenStudy (amistre64):

simple or compunded?

OpenStudy (amistre64):

i know i hit the "o" ... c o m p o u n d e d.

OpenStudy (anonymous):

simple.

OpenStudy (amistre64):

Prt then 300(.05)(7)

OpenStudy (amistre64):

and mathically, you can move the decimal around if by hand is tough 300(.05)(7) = 3(5)(7)

OpenStudy (anonymous):

\[S.I = \frac{ P \times R \times T}{ 100 }\] S.I=Simple Interest P=Principal R=Rate T=Time